# SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining

### v3

@@ Abstract
The quadratic complexity of self-attention in Transformer-based Large Language Models (LLMs) renders long-context inference prohibitively expensive. While Sliding Window Attention (SWA), the simplest sparse attention pattern, offers a linear-complexity alternative, naively applying it to models pretrained with Full Attention (FA) causes catastrophic long-context performance collapse due to the training–inference mismatch. To address this, we propose Sliding Window Attention Adaptation (SWAA), a plug-and-play toolkit of recipes that adapt FA models to SWA without costly pretraining. SWAA systematically combines five strategies: (1) applying SWA only during prefilling; (2) preserving “sink” tokens; (3) interleaving FA/SWA layers; (4) chain-of-thought (CoT); and (5) fine-tuning. Our experiments demonstrate that while individual methods are insufficient, specific synergistic combinations can effectively recover original long-context capabilities. After further analyzing performance-efficiency trade-offs, we identify recommended SWAA configurations for diverse scenarios, which achieve 30% to 100% speedups for long-context LLM inference with acceptable quality loss. Our code is available at github.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) vaswani2017attention demonstrate remarkable capabilities, but their self-attention scales quadratically with the input sequence length, making long context processing inefficient. Sliding Window Attention (SWA), the most straightforward and widely adopted sparse attention pattern, which restricts each token’s attention to a fixed-size local window, reduces the computational complexity to linearity, along with some other benefits (see Appendix A).
Despite its simplicity and ease of use, significant challenges remain in leveraging SWA to accelerate LLM inference. On one hand, due the performance degradation brought by training-inference attention pattern mismatch, to apply SWA to LLMs, most typical solutions involve training a model with SWA from scratch, but they are prohibitively costly and cannot match the performance of state-of-the-art full-causal-attention models like Qwen3 yang2025qwen3technicalreport, mainly due to the inability to reproduce pretraining data. On the other hand, training-free methods like streaming attention xiao2024efficientstreaminglanguagemodels can stabilize LLM outputs by retaining “sink tokens” while applying SWA, which greatly improve efficiency, but inevitably suffer from severe long-context performance degradation possibly due to the inaccessibility of distant tokens’ information xiao2025sliding. This motivates a critical, unexplored question: Can we adapt a full-attention model to sliding window attention at low cost while maintaining long-context performance?
@@
Among these, FA Decode is a novel method we introduce. Keep First $k$ Tokens and FA/SWA Interleaving have been proven effective in prior work xiao2024efficientstreaminglanguagemodels; gemmateam2024gemma2improvingopen; zhang2025lighttransferlongcontextllmsecretly, while CoT and fine-tuning are common LLM techniques. However, the optimal strategy for combining these methods to achieve effective SWA adaptation remains unexplored.
Therefore, in our experiments, we evaluate SWAA on Qwen3 yang2025qwen3technicalreport and Llama3.1 grattafiori2024llama3herdmodels across several long-context benchmarks, measuring both performance and efficiency under a wide range of SWAA configurations. First, we find that each method makes a distinct contribution, but no single ingredient suffices to make SWA competitive with full attention. Second, we show that specific synergistic combinations of methods can recover a large fraction of the original long-context performance. Third, we analyze the performance-efficiency trade-offs of different SWAA recipes and identify some recommended configurations suitable for different deployment scenarios, which can achieving nearly 100% speedup with 90% accuracy retention in efficiency-first scenarios, or 30% speedup with nearly 100% accuracy retention in quality-first scenarios. This shows SWAA is a flexible toolkit beyond a fixed strategy.
@@
    We implement our methods with Flash-Attention dao2023flashattention2fasterattentionbetter and vLLM kwon2023efficientmemorymanagementlarge, making it plug-and-play and user-friendly for practical deployment.
@@ 2 Related Works
The $O(N^{2})$ complexity of self-attention in Transformers vaswani2017attention has spurred a wide field of research about more efficient language model architectures. Among the two most popular technological routes are sparse attention and linear attention.
@@ 2.1 Sparse Attention
Our work falls in this category. Sliding Window Attention (SWA) represents the most basic form of local sparse attention, yet its performance is inherently limited. Therefore, model architectures such as Longformer beltagy2020longformerlongdocumenttransformer, BigBird zaheer2021bigbirdtransformerslonger, and RATTENTION wang2025rattentionminimalslidingwindow combine local SWA on most tokens with special global attention on specific tokens to create a more powerful, albeit still sparse, pattern. Popular LLMs like Gemma2 gemmateam2024gemma2improvingopen adopt SWA in half of their layers to balance the efficiency of SWA and performance of FA. Sliding Window Attention Training (SWAT) fu2025slidingwindowattentiontraining introduces architectural changes, such as sigmoid activation and balanced position embeddings, to stabilize SWA performance. More advanced methods like Deepseek-sparse-attention yuan-etal-2025-native; deepseekai2025deepseekv32pushingfrontieropen, although achieving excellent quality, involve more complicated implementation and optimization due to semantic-aware attention operations (e.g., selecting the most important tokens based on attention weights). Nevertheless, nearly all aforementioned methods require pretraining with a specific sparse pattern, which is costly and fails to leverage the advantages of existing pretrained models.
LightTransfer zhang2025lighttransferlongcontextllmsecretly is a promising attempt at adapting existing models to SWA without pretraining, which has the same motivation as ours. But it may generalize poorly across model families (see Appendix F).
@@ 2.2 Linear Attention
An alternative approach involves reformulating the attention mechanism entirely to achieve linear, $O(N)$, complexity. This includes methods such as RNN-like linear attention transformers katharopoulos_et_al_2020; peng2023rwkvreinventingrnnstransformer; sun2023retentivenetworksuccessortransformer and structured state-space models (SSMs) like Mamba gu2024mambalineartimesequencemodeling. Many works such as Jamba and Nemotron-Flashlieber2024jambahybridtransformermambalanguage; noauthor_bamba_nodate; lingteam2025attentionmattersefficienthybrid; funemotron interleave linear attention layers with traditional attention layers to create hybrid model structures. While promising, they represent a fundamental architectural departure from the standard Transformer, and their performance is generally weaker than traditional Transformer-based LLMs.
@@ 3.1.1 Keep First k Tokens
Streaming attention xiao2024efficientstreaminglanguagemodels demonstrates that models pretrained with full attention allocate a disproportionate amount of attention to the initial tokens ("attention sink"), so it permanently retains the attention to these "sink" tokens while using SWA, which successfully maintains the stability of the attention distribution and the model’s output. Our method is basically the same: as shown in Figure 1(b), any subsequent token can attend to its local window and the initial $k$ tokens.
Crucially, we extend this approach further. Streaming attention operates only at the KV cache level; specifically, the KV cache of the sink tokens is retained during decoding, while the prefilling stage is not modified or accelerated at all. In contrast, we directly customize the Flash-Attention-2 dao2023flashattention2fasterattentionbetter kernel to implement such attention mask, thereby accelerating prefilling via SWA as well, and eliminating the need to modify KV cache.
@@ 3.2 Interleaving Layers
This method retains full attention on a subset of layers while applying SWA to the remainder, providing a simple hybrid mechanism to balance the performance of full attention with the efficiency of pure SWA. A common strategy involves designating one in every $n$ layers to use full attention (e.g., layers 0, 2, 4, …retain full attention, while all others use SWA). For example, Gemma-2 gemmateam2024gemma2improvingopen uses SWA only for layers [1, 3, 5, …], and Gemma-3 gemmateam2025gemma3technicalreport uses SWA only for layers [5, 11, 17, …].
However, for an FA-pretrained model, layers may exhibit distinct behaviors, suggesting it may not be optimal to simply assign SWA to layers [1, 3, 5, …]. Some methods like LightTransfer zhang2025lighttransferlongcontextllmsecretly have explored more fine-grained layer selection, but we find it is not consistently superior in practice (see Appendix F). Therefore, for simplicity, we still choose the simplest fixed-interval layer selections in our experiments, such as [0, 2, 4, …] and [1, 3, 5, …].
@@ 3.3 Chain-of-Thought
Chain-of-Thought (CoT) wei2023chainofthoughtpromptingelicitsreasoning is a widely used technique for improving model accuracy via reasoning. With the advent of "thinking" models, such as DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability, CoT has evolved from a prompting strategy to an intrinsic LLM capability. However, whether CoT has a specific impact in SWA scenarios remains uninvestigated. We explore this by comparing a thinking model with its non-thinking variant, e.g., Qwen3-4B-Thinking and Qwen3-4B-Instruct, to verify the effect of CoT on SWA adaptation, which can produce more notable differences compared to simply adding CoT prompting to the same model.
@@ 3.4 Fine-tuning
However, most available long-context datasets only contain brief ground-truth answers, lacking the reasoning steps required for fine-tuning a "thinking" model. Since our goal is to restore the model’s original capabilities under SWA rather than teach it new ones, we adopt an approach similar to self-distillation yang2024selfdistillationbridgesdistributiongap. Specifically, we utilize the original full-attention model to generate new answers for the dataset’s questions, and these generated answers are then filtered for correctness using GPT-5-Mini noauthor_chatgpt_nodate, to construct our training dataset. For each question, we sample 4 answers with temperature 1 to enhance diversity, because we find it is slightly better than generating only one answer.
@@
19 2k [1, 5, 9, …] 10 True 53.2 31.4 32.4 26.4 9.8 26.4
20 2k [1, 9, 17, …] 10 True 36.4 18.8 30.5 25.4 0.6 0.2
21 2k [3, 7, 11, …] 10 True 54.2 34.6 29.8 26.0 1.6 4.2
22 4k [1, 3, 5, …] 100 True 73.0 54.2 35.6 31.2 65.4 83.6
23 8k [1, 3, 5, …] 100 True 71.6 56.6 35.8 31.5 67.2 86.0
@@ 4.1 Models
Our primary experiments use Qwen3-4B-Thinking and Qwen3-4B-Instruct yang2025qwen3technicalreport. The Thinking variant enforces chain-of-thought (CoT) style reasoning, whereas the Instruct variant usually just answers briefly. To ensure generality, we additionally evaluate Qwen3-30B-A3B-Thinking, Qwen3-30B-A3B-Instruct yang2025qwen3technicalreport, and Llama3.1-8B-Instruct touvron2023llamaopenefficientfoundation, as shown in Appendix C.
All models are served with vLLM in float16 precision using a batch size of 64. We use greedy decoding (temperature =0=0) for all evaluations. In preliminary experiments, we observed that vLLM yields slightly lower (about 1% to 5%) scores than HuggingFace Transformers due to precision-related discrepancies.
@@ 4.2 Evaluation Dataset
SWA is identical to full attention when the context length is within the window size. Even if the model is fine-tuned, we can simply disable the LoRA adapters for short prompts to get the same response as the original model. Therefore, our evaluations focus on long-context benchmarks, as re-evaluating models on standard short-context benchmarks (e.g., MMLU hendrycks2021measuringmassivemultitasklanguage, GPQA rein2023gpqagraduatelevelgoogleproofqa) is completely unnecessary.
We select LongMemEval wu2025longmemevalbenchmarkingchatassistants, a benchmark consisting of various types of long-context QA tasks with moderate difficulty, as our main benchmark, although it is originally designed for agent memory system evaluation. Its context length is controllable by selecting a specific number of chat sessions to concatenate as the context from a pool of hundreds of sessions (a session contains the chat history between user and assistant within a day). To create a moderately difficult and discriminative evaluation, we construct LongMemEval_24k by sampling 10 sessions, resulting in 500 samples ranging from 16k to 32k with an average context length of 24k.
We also experiment on LongBench-V2 bai2025longbenchv2deeperunderstanding and Ruler hsieh2024rulerwhatsrealcontext, which are most commonly used long-context benchmarks, for additional validation of generalizability. For LongBench V2, we retain only the samples whose context length is under 128k due to GPU memory limitations; thus, 311 of 500 samples are kept. For Ruler, we choose the Multi-Query task, which contains 500 samples, and control the context length to 128k (counted by Qwen3 tokenizer).
To judge the correctness for accuracy, we use LLM-as-judge with GPT-5-Mini noauthor_chatgpt_nodate for LongMemEval and exact-match for LongBench-V2 and Ruler.
@@ 4.3 Training Details
For the fine-tuning dataset, we initially considered LongAlign bai2024longalignrecipelongcontext, a widely used long-context fine-tuning dataset for adapt a regular-length model to long-context tasks. However, since its sample count ($\sim$10,000) is insufficient, we incorporate an additional 6,000 samples from Fusang-v1-long Fusang-V1, a more comprehensive corpus of over 40,000 long-context samples that includes LongAlign as a subset.
We perform SWA-aware fine-tuning using LoRA hu2021loralowrankadaptationlarge. Unless otherwise noted, we use rank $r=16$ and $\alpha=128$, and apply LoRA only to the query, key, and value projection modules. We adopt this parameter-efficient setting because full-parameter fine-tuning often leads to overfitting and degradation of the model’s original capabilities in our preliminary experiments. We use a learning rate of 1e-4 with a cosine decay schedule. Models are fine-tuned for a single epoch on the sampled long-context dataset since we observe no meaningful gains from additional epochs (see Appendix E). Training of each SWAA configuration takes approximately 12 hours on an 8*H20 GPU server for Qwen3-4B and 30 hours for Qwen3-30B-A3B.

### v4

@@ Abstract
The quadratic complexity of self-attention in Transformer-based Large Language Models (LLMs) renders long-context inference prohibitively expensive. While Sliding Window Attention (SWA), the simplest sparse attention pattern, offers a linear-complexity alternative, naively applying it to models pretrained with Full Attention (FA) causes catastrophic long-context performance collapse due to the training-inference mismatch. To address this, we propose Sliding Window Attention Adaptation (SWAA), a plug-and-play toolkit of recipes that adapt FA models to SWA without costly pretraining. SWAA systematically combines five strategies: (1) applying SWA only during prefilling; (2) preserving “sink” tokens; (3) interleaving FA/SWA layers; (4) chain-of-thought (CoT); and (5) fine-tuning. Our experiments demonstrate that while individual methods are insufficient, specific synergistic combinations can effectively recover original long-context capabilities. After further analyzing performance-efficiency trade-offs, we identify recommended SWAA configurations for diverse scenarios, which achieve 30% to 100% speedups for long-context LLM inference with acceptable quality loss. Our code is available at github.
@@ 1 Introduction
Transformer-based Large Language Models (LLMs) Vaswani et al. (2017) demonstrate remarkable capabilities, but their self-attention scales quadratically with the input sequence length, making long context processing inefficient. Sliding Window Attention (SWA), the most straightforward and widely adopted sparse attention pattern, which restricts each token’s attention to a fixed-size local window, reduces the computational complexity to linearity, along with some other benefits (see Appendix A).
Despite its simplicity and ease of use, significant challenges remain in leveraging SWA to accelerate LLM inference. On one hand, due the performance degradation brought by training-inference attention pattern mismatch, to apply SWA to LLMs, most typical solutions involve training a model with SWA from scratch, but they are prohibitively costly and cannot match the performance of state-of-the-art full-causal-attention models like Qwen3 Team (2025b), mainly due to the inability to reproduce pretraining data. On the other hand, training-free methods like streaming attention Xiao et al. (2024) can stabilize LLM outputs by retaining “sink tokens” while applying SWA, which greatly improve efficiency, but inevitably suffer from severe long-context performance degradation possibly due to the inaccessibility of distant tokens’ information Xiao (2025). This motivates a critical, unexplored question: Can we adapt a full-attention model to sliding window attention at low cost while maintaining long-context performance?
@@
Among these, FA Decode is a novel method we introduce. Keep First $k$ Tokens and FA/SWA Interleaving have been proven effective in prior work Xiao et al. (2024); Team (2024a); Zhang et al. (2024), while CoT and fine-tuning are common LLM techniques. However, the optimal strategy for combining these methods to achieve effective SWA adaptation remains unexplored.
Therefore, in our experiments, we evaluate SWAA on Qwen3 Team (2025b) and Llama3.1 Team (2024b) across several long-context benchmarks, measuring both performance and efficiency under a wide range of SWAA configurations. First, we find that each method makes a distinct contribution, but no single ingredient suffices to make SWA competitive with full attention. Second, we show that specific synergistic combinations of methods can recover a large fraction of the original long-context performance. Third, we analyze the performance-efficiency trade-offs of different SWAA recipes and identify some recommended configurations suitable for different deployment scenarios, which can achieving nearly 100% speedup with 90% accuracy retention in efficiency-first scenarios, or 30% speedup with nearly 100% accuracy retention in quality-first scenarios. This shows SWAA is a flexible toolkit beyond a fixed strategy.
@@
    We implement our methods with Flash-Attention Dao (2024) and vLLM Kwon et al. (2023), making it plug-and-play and user-friendly for practical deployment.
@@ 2 Related Works
The $O(N^{2})$ complexity of self-attention in Transformers Vaswani et al. (2017) has spurred a wide field of research about more efficient language model architectures. Among the two most popular technological routes are sparse attention and linear attention.
@@ 2.1 Sparse Attention
Our work falls in this category. Sliding Window Attention (SWA) represents the most basic form of local sparse attention, yet its performance is inherently limited. Therefore, model architectures such as Longformer Beltagy et al. (2020), BigBird Zaheer et al. (2020), and RATTENTION Wang et al. (2025) combine local SWA on most tokens with special global attention on specific tokens to create a more powerful, albeit still sparse, pattern. Popular LLMs like Gemma2 Team (2024a) adopt SWA in half of their layers to balance the efficiency of SWA and performance of FA. Sliding Window Attention Training (SWAT) Fu et al. (2025b) introduces architectural changes, such as sigmoid activation and balanced position embeddings, to stabilize SWA performance. More advanced methods like Deepseek-sparse-attention Yuan et al. (2025); DeepSeek-AI (2025b), although achieving excellent quality, involve more complicated implementation and optimization due to semantic-aware attention operations (e.g., selecting the most important tokens based on attention weights). Nevertheless, nearly all aforementioned methods require pretraining with a specific sparse pattern, which is costly and fails to leverage the advantages of existing pretrained models.
LightTransfer Zhang et al. (2024) is a promising attempt at adapting existing models to SWA without pretraining, which has the same motivation as ours. But it may generalize poorly across model families (see Appendix F).
@@ 2.2 Linear Attention
An alternative approach involves reformulating the attention mechanism entirely to achieve linear, $O(N)$, complexity. This includes methods such as RNN-like linear attention transformers Katharopoulos et al. (2020); Peng et al. (2023); Sun et al. (2023) and structured state-space models (SSMs) like Mamba Gu and Dao (2023). Many works such as Jamba and Nemotron-FlashLieber et al. (2024); Linsong Chu and others (2024); Team et al. (2025); Fu et al. (2025a) interleave linear attention layers with traditional attention layers to create hybrid model structures. While promising, they represent a fundamental architectural departure from the standard Transformer, and their performance is generally weaker than traditional Transformer-based LLMs.
@@ 3.1.1 Keep First k Tokens
Streaming attention Xiao et al. (2024) demonstrates that models pretrained with full attention allocate a disproportionate amount of attention to the initial tokens ("attention sink"), so it permanently retains the attention to these "sink" tokens while using SWA, which successfully maintains the stability of the attention distribution and the model’s output. Our method is basically the same: as shown in Figure 1(b), any subsequent token can attend to its local window and the initial $k$ tokens.
Crucially, we extend this approach further. Streaming attention operates only at the KV cache level; specifically, the KV cache of the sink tokens is retained during decoding, while the prefilling stage is not modified or accelerated at all. In contrast, we directly customize the Flash-Attention-2 Dao (2024) kernel to implement such attention mask, thereby accelerating prefilling via SWA as well, and eliminating the need to modify KV cache.
@@ 3.2 Interleaving Layers
This method retains full attention on a subset of layers while applying SWA to the remainder, providing a simple hybrid mechanism to balance the performance of full attention with the efficiency of pure SWA. A common strategy involves designating one in every $n$ layers to use full attention (e.g., layers 0, 2, 4, …retain full attention, while all others use SWA). For example, Gemma-2 Team (2024a) uses SWA only for layers [1, 3, 5, …], and Gemma-3 Team (2025a) uses SWA only for layers [5, 11, 17, …].
However, for an FA-pretrained model, layers may exhibit distinct behaviors, suggesting it may not be optimal to simply assign SWA to layers [1, 3, 5, …]. Some methods like LightTransfer Zhang et al. (2024) have explored more fine-grained layer selection, but we find it is not consistently superior in practice (see Appendix F). Therefore, for simplicity, we still choose the simplest fixed-interval layer selections in our experiments, such as [0, 2, 4, …] and [1, 3, 5, …].
@@ 3.3 Chain-of-Thought
Chain-of-Thought (CoT) Wei et al. (2022) is a widely used technique for improving model accuracy via reasoning. With the advent of "thinking" models, such as DeepSeek-R1 DeepSeek-AI (2025a), CoT has evolved from a prompting strategy to an intrinsic LLM capability. However, whether CoT has a specific impact in SWA scenarios remains uninvestigated. We explore this by comparing a thinking model with its non-thinking variant, e.g., Qwen3-4B-Thinking and Qwen3-4B-Instruct, to verify the effect of CoT on SWA adaptation, which can produce more notable differences compared to simply adding CoT prompting to the same model.
@@ 3.4 Fine-tuning
However, most available long-context datasets only contain brief ground-truth answers, lacking the reasoning steps required for fine-tuning a "thinking" model. Since our goal is to restore the model’s original capabilities under SWA rather than teach it new ones, we adopt an approach similar to self-distillation Yang et al. (2024). Specifically, we utilize the original full-attention model to generate new answers for the dataset’s questions, and these generated answers are then filtered for correctness using GPT-5-Mini OpenAI (2025), to construct our training dataset. For each question, we sample 4 answers with temperature 1 to enhance diversity, because we find it is slightly better than generating only one answer.
@@
19 2k [1, 5, 9, …] 10 True 53.2 31.4 29.6 26.4 9.8 26.4
20 2k [1, 9, 17, …] 10 True 36.4 18.8 28.6 25.4 0.6 0.2
21 2k [3, 7, 11, …] 10 True 54.2 34.6 27.7 26.0 1.6 4.2
22 4k [1, 3, 5, …] 100 True 73.0 54.2 33.4 31.2 65.4 83.6
23 8k [1, 3, 5, …] 100 True 71.6 56.6 35.4 31.5 67.2 86.0
@@ 4.1 Models
Our primary experiments use Qwen3-4B-Thinking and Qwen3-4B-Instruct Team (2025b). The Thinking variant enforces chain-of-thought (CoT) style reasoning, whereas the Instruct variant usually just answers briefly. To ensure generality, we additionally evaluate Qwen3-30B-A3B-Thinking, Qwen3-30B-A3B-Instruct Team (2025b), and Llama3.1-8B-Instruct Touvron et al. (2023), as shown in Appendix C.
All models are served with vLLM in float16 precision using a batch size of 64, with our customized Flash-attention-2 kernel to support Keep First and FA Decoding. We use greedy decoding (temperature =0=0) for all evaluations. In preliminary experiments, we observed that vLLM yields slightly lower (about 1% to 5%) scores than HuggingFace Transformers due to precision-related discrepancies.
@@ 4.2 Evaluation Dataset
SWA is identical to full attention when the context length is within the window size. Even if the model is fine-tuned, we can simply disable the LoRA adapters for short prompts to get the same response as the original model. Therefore, our evaluations focus on long-context benchmarks, as re-evaluating models on standard short-context benchmarks (e.g., MMLU Hendrycks et al. (2021), GPQA Rein et al. (2023)) is completely unnecessary.
We select LongMemEval Wu et al. (2024), a benchmark consisting of various types of long-context QA tasks with moderate difficulty, as our main benchmark, although it is originally designed for agent memory system evaluation. Its context length is controllable by selecting a specific number of chat sessions to concatenate as the context from a pool of hundreds of sessions (a session contains the chat history between user and assistant within a day). To create a moderately difficult and discriminative evaluation, we construct LongMemEval_24k by sampling 10 sessions, resulting in 500 samples ranging from 16k to 32k with an average context length of 24k.
We also experiment on LongBench-V2 Bai et al. (2024b) and Ruler Hsieh et al. (2024), which are most commonly used long-context benchmarks, for additional validation of generalizability. For LongBench V2, we retain only the samples whose context length is under 128k due to GPU memory limitations; thus, 311 of 500 samples are kept. For Ruler, we choose the Multi-Query task, which contains 500 samples, and control the context length to 128k (counted by Qwen3 tokenizer).
To judge the correctness for accuracy, we use LLM-as-judge with GPT-5-Mini OpenAI (2025) for LongMemEval and exact-match for LongBench-V2 and Ruler.
@@ 4.3 Training Details
For the fine-tuning dataset, we initially considered LongAlign Bai et al. (2024a), a widely used long-context fine-tuning dataset for adapt a regular-length model to long-context tasks. However, since its sample count ($\sim$10,000) is insufficient, we incorporate an additional 6,000 samples from Fusang-v1-long Pan (2024), a more comprehensive corpus of over 40,000 long-context samples that includes LongAlign as a subset.
We perform SWA-aware fine-tuning using LoRA Hu et al. (2022). Unless otherwise noted, we use rank $r=16$ and $\alpha=128$, and apply LoRA only to the query, key, and value projection modules. We adopt this parameter-efficient setting because full-parameter fine-tuning often leads to overfitting and degradation of the model’s original capabilities in our preliminary experiments. We use a learning rate of 1e-4 with a cosine decay schedule. Models are fine-tuned for a single epoch on the sampled long-context dataset since we observe no meaningful gains from additional epochs (see Appendix E). Training of each SWAA configuration takes approximately 12 hours on an 8*H20 GPU server for Qwen3-4B and 30 hours for Qwen3-30B-A3B.